Pentagon has a perimeter of 9/10 meter. The equation 5s=9/10represents the lengah each side. How is s solved
Answer by macston(5194)

5s=9/10m Multiply each side by 1/5
s=9/50m=0.18m ANSWER: Each side is 0.18 meter.